Show that the sequence (fn) sequence where


fn(x)= x/(1+nx^2), x∈[2,∞] is uniformly convergent in [2,∞]

1
Expert's answer
2022-02-09T12:08:09-0500

ANSWER .

Let "x" be fixed and "x\in [2,\infty )" . Then "f(x)=\lim _{ n\rightarrow \infty }{ { f }_{ n } } \left( x \right) =\lim _{ n\rightarrow \infty }{ \frac { x }{ 1+n{ x }^{ 2 } } = } x\lim _{ n\rightarrow \infty }{ \left( \frac { 1 }{ n } \cdot \frac { 1 }{ \frac { 1 }{ n } +{ x }^{ 2 } } \right) = }" "=x\left( \lim _{ n\rightarrow \infty }{ \left( \frac { 1 }{ n } \ \right) \cdot \frac { 1 }{ { x }^{ 2 }+\lim _{ n\rightarrow \infty }{ \frac { 1 }{ n } } } } \right) =0" . So "f_n" converges pointwise to "f(x)=0."

Denote "{ d }_{ n }=\underset { 2\le x<\infty }{ sup } \left| { f }_{ n }\left( x \right) -f\left( x \right) \right|" . Since "{ f }_{ n }^{ ' }\left( x \right) =\frac { 1+n{ x }^{ 2 }-2n{ x }^{ 2 } }{ { \left( 1+n{ x }^{ 2 } \right) }^{ 2 } } =\frac { 1-n{ x }^{ 2 } }{ { \left( 1+n{ x }^{ 2 } \right) }^{ 2 } } <0^{}" for all "x\ge 2" "\left( n\ge 1 \right)" ,then each function "f_n" decreases on the interval "[2,\infty )"

Therefore, "\underset { 2\le x<\infty }{ sup } \frac { x }{ 1+n{ x }^{ 2 } } ={ f }_{ n }\left( 2 \right) =\frac { 2 }{ 1+4n }" . So, "{ d }_{ n }=\underset { 2\le x<\infty }{ sup } \left| \frac { x }{ 1+n{ x }^{ 2 } } -0 \right| =\frac { 2 }{ 1+4n }" . Since , "\lim _{ n\rightarrow \infty }{ { d }_{ n }= } \lim _{ n\rightarrow \infty }{ \frac { 2 }{ 1+4n } =0 }", then by definition, we get that the sequence converges uniformly in "[2,\infty )" .